An ice rink is in the shape of a rectangle with semicircular ends. The straight
sides are 60m long and 30m apart. It is to be filled to a depth of 10cm with
ice. Calculate how many cubic metres of ice will be needed.

Found 2 solutions by fractalier, Boreal:
Answer by fractalier(6550)

The volume of ice needed is the area of the rink times the depth, 0.1 m.
The rink is a 60m by 30m rectangle. That area is 1800 m^2.
The semi-circles at each end become a full circle whose radius is 15m
Its area is A = (pi)r^2 = (3.14)(15^2) = 706.5 m^2
Thus the total area of the rink is
1800 + 706.5 = 2506.5 m
Now multiply that by the depth (0.1m) and we get
V = 2506.5 * 0.1 = 250.65 cubic meters of ice.

Answer by Boreal(15235)

It helps to draw this.
There is a rectangle with dimensions 60 m X 30 m.
The two semicircular ends form one circle with diameter 30 m, radius 15 m.
The area is 1800 m^2+225*pi m^2, the last being the area of a circle.
The total area is 2506.85 m^2, after multiplying by pi.
This is to be multiplied by 10 cm=0.1 m
The units will be m^3.
250.86 m^3 of ice are needed after multiplying by 0.1